\begin{thebibliography}{99}
\bibitem{Lucha07:HORSE}Z.-F.~Li, W.~Lucha, and F.~F.~Sch\"oberl,
\emph{Phys.~Rev.~D\/} \textbf{76}, 125028 (2007) [arXiv:0707.3202
[hep-ph]].
\bibitem{Lucha07:StabOSS-QCD@Work07}W.~Lucha and F.~F.~Sch\"oberl,
in \emph{QCD@Work 2007 --- International Workshop on Quantum
Chromodynamics --- Theory and Experiment\/}, edited by
P.~Colangelo et al., AIP Conf.~Proc., AIP, Melville, New York,
2007, Vol.~\textbf{964}, p.~318 [arXiv:0707.1440 [hep-ph]].
\bibitem{Lucha05:IBSEWEP(a)}W.~Lucha and F.~F.~Sch\"oberl,
\emph{J.~Phys.~G: Nucl.~Part.~Phys.\/} \textbf{31}, 1133 (2005)
[hep-th/0507281].
\bibitem{Lucha05:IBSEWEP(b)}W.~Lucha and F.~F.~Sch\"oberl, in
\emph{Quark Confinement and the Hadron Spectrum VII --- 7\/$^{\it
th}$ Conference on Quark Confinement and the Hadron Spectrum ---
QCHS7\/}, edited by J.~E.~F.~T.~Ribeiro, AIP Conf.~Proc., AIP,
Melville, New York, 2007, Vol.~\textbf{892}, p.~524
[hep-ph/0610016].
\bibitem{Lucha05:EQPIBSE}Z.-F.~Li, W.~Lucha, and F.~F.~Sch\"oberl,
\emph{Mod.~Phys.~Lett.~A\/} \textbf{21}, 1657 (2006)
[hep-ph/0510372].
\bibitem{Lucha07:SSSECI-Hadron07}W.~Lucha and F.~F.~Sch\"oberl, in
\emph{XII International Conference on Hadron Spectroscopy ---
Hadron~07\/}, edited by L.~Benussi et al., Frascati Phys.~Ser.,
INFN, Laboratori Nazionali di Frascati, 2007, Vol.~\textbf{46},
p.~1539 [arXiv:0711.1736~[hep-ph]].
\bibitem{Lucha07:HORSEWEP}Z.-F.~Li, W.~Lucha, and F.~F.~Sch\"oberl,
\emph{J.~Phys.~G: Nucl.~Part.~Phys.\/} \textbf{35}, 115002 (2008)
[arXiv:0712.2947 [hep-ph]].
\bibitem{Lucha08-C8}W.~Lucha, in \emph{8\/$^{\it th}$ International
Conference on Quark Confinement and the Hadron Spectrum --- QCHS
2008\/}, PoS (Confinement8) 164 (2009) [arXiv:0810.5500 [hep-ph]].
\bibitem{MMP-OI(a)}W.~Lucha and F.~F.~Sch\"oberl,
\emph{Int.~J.~Mod.~Phys.~A\/} \textbf{14},~2309 (1999)
[hep-ph/9812368].
\bibitem{MMP-OI(b)}W.~Lucha and F.~F.~Sch\"oberl, \emph{Fizika B\/}
\textbf{8}, 193 (1999) [hep-ph/9812526].
\bibitem{MMP-OI(c)}W.~Lucha and F.~F.~Sch\"oberl,
\emph{Phys.~Rev.~A\/} \textbf{54}, 3790 (1996) [hep-ph/9603429].
\bibitem{MMP-OI(d)}W.~Lucha and F.~F.~Sch\"oberl,
\emph{J.~Math.~Phys.\/} \textbf{41}, 1778 (2000) [hep-ph/9905556].
\bibitem{MMP-OI(e)}W.~Lucha and F.~F.~Sch\"oberl, \emph{Recent
Res.~Devel.~Physics\/}~\textbf{5}, 1423 (2004) [hep-ph/0408184].
\bibitem{MMP-OI(f)}Z.-F.~Li, J.-J.~Liu, W.~Lucha, W.-G.~Ma, and
F.~F.~Sch\"oberl, \emph{J.~Math.~Phys.\/} \textbf{46}, 103514
(2005) [hep-ph/0501268].
\bibitem{Lucha10:QCD@Work10}W.~Lucha, arXiv:1008.1404 [hep-ph], to
be published in the Proceedings of \emph{QCD@Work 2010 ---
5\/$^{\it th}$ International Workshop on Quantum Chromodynamics
--- Theory and Experiment\/}, AIP Conf.~Proc., AIP, Melville, New
York.
\bibitem{Lucha:IntEq}W.~Lucha and F.~F.~Sch\"oberl (in
preparation).
\end{thebibliography}

\documentclass[sort&compress,final]{aipproc}\layoutstyle{8x11double}
\begin{document}\title{Analytic Quest for Confining Interaction
Kernels in Instantaneous Bethe--Salpeter Equations}
\classification{11.10.St, 03.65.Pm, 03.65.Ge, 12.38.Aw}
\keywords{Bethe--Salpeter formalism, three-dimensional reduction,
instantaneous approximation, generalized Salpeter equation,
description of bound states within quantum field theories,
spectral analysis, confining interactions, Lorentz structures of
Bethe--Salpeter interaction kernels}\author{Wolfgang
LUCHA}{address={Institute for High Energy Physics, Austrian
Academy of Sciences, Nikolsdorfergasse 18, A-1050 Vienna,
Austria}}\maketitle

\noindent The Salpeter equation is a well-known three-dimensional
approximation to the Bethe--Salpeter equation crucial for the
description of bound states by quantum field theories. Therefore,
it represents a standard tool in hadron physics. It is obtained by
assuming the bound-state constituents to propagate as free
particles (with effective masses) and by considering their
interactions in the `instantaneous limit'.

However, depending on the specific Lorentz behaviour of the
Bethe--Salpeter interaction kernel, implementation of confinement
in the Salpeter equation in too na\"ive ways might lead to the
prediction of unstable states where only truly bound states are
expected: Their energy eigenvalues may be embedded in a continuous
spectrum, for instance.

These observations call for a rigorous spectral analysis of the
Salpeter equation, in order to pin down its essential spectral
features by regarding a bound state as stable if its energy (or
mass) eigenvalue belongs to a real and discrete part of the
spectrum that is bounded from below. Reusing ideas and methods
designed in our previous studies of the `reduced Salpeter
equation' \cite{Lucha07:HORSE,Lucha07:StabOSS-QCD@Work07} and of
its improvement \cite{Lucha05:IBSEWEP(a),Lucha05:IBSEWEP(b),
Lucha05:EQPIBSE} derived by allowing for `dressed' propagators of
all bound-state constituents \cite{Lucha07:SSSECI-Hadron07,
Lucha07:HORSEWEP}, we therefore embark on a thorough investigation
of full --- in contrast to reduced --- Salpeter equations for
fermion--antifermion bound states.

We exploit the serendipity that describing confinement by a
configuration-space potential of harmonic-oscillator form reduces
the Salpeter integral equation to a system of (radial) eigenvalue
differential equations. The analysis of confining interactions of
time-component Lorentz-vector nature is comparatively easy and the
stability of all bound states may be established analytically
\cite{Lucha08-C8} by constructing operator inequalities and
applying the characterization of all discrete eigenvalues of
self-adjoint operators bounded from below by the minimum-maximum
principle \cite{MMP-OI(a),MMP-OI(b),MMP-OI(c),MMP-OI(d),MMP-OI(e),
MMP-OI(f)}. Regrettably, the proof cannot be transferred to the
case of a confining interaction of Lorentz-scalar or -pseudoscalar
structure but for a linear combination of time-component
Lorentz-vector and Lorentz-scalar confining interactions stability
is assured (irrespective of the relative sign of our two
contributions) if the vector dominates the scalar
\cite{Lucha10:QCD@Work10}. `Non-harmonically', we will face
integral equations \cite{Lucha:IntEq}.

\bibliographystyle{aipproc}\end{document}